you invested money in two funds. last year, the first fund paid a dividend of 9% and the second a dividend of 3, and you received a total of $900. this year, the first fund paid a 10% dividend and the second only 1%, and you received a total of $860. how much money did you invest in each total fund?
Answer by mananth(16946)

x Investment in Fund I
y Investment in Fund II
Interest calculation in the two repective years.
0.09x+0.03y= 900...............(1)
0.10x+0.01y =860...............(2)
Multiply both equations by 100 to eliminate decimals
9 x + 3 y = 90000 .............1

10 x + 1 y = 86000 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -3
9 x 3 y = 90000
-30 x -3 y = -258000
Add the two equations
-21 x = -168000
/ -21
x = 8000
plug value of x in (1)
9 x + 3 y = 90000
72000 + 3 y = 90000
3 y = 90000-72000
3 y = 18000
y = 6000
$ 8000 Investment in Fund I
$ 6000 Investment in Fund II
m.ananth@hotmail.ca